# Learn while Unlearn: An Iterative Unlearning Framework for Generative Language Models

### v2

@@ 4.4 Configurations
4.5 Main results
The results of all methods are summarized in Table 1. Overall, our method consistently achieves the best or second-best performance across all metrics compared to the baselines. Although methods like KUMPR, DPO, KL and LLMU may exhibit seemingly better unlearning ability measured by EL, MA, and BERT, our method preserves the generation ability of the model. Compared to Opt, our model unlearns the original model better. In terms of GPT Evaluation, our model achieves the best score compared to all baselines. Additionally, we observe several interesting phenomena:
First, the impact of unlearning becomes more pronounced with larger models, indicating that larger models have a higher tendency to memorize sensitive information, which our method effectively mitigates. Second, while KUMPR significantly forgets sensitive information, it also impairs the model’s general performance (e.g., PPL). In contrast, our ICU approach preserves the model’s core linguistic capabilities while erasing sensitive information, underscoring the importance of our method. Full results for each sample are provided in Appendix G.
4.6 Ablation study
4.7 Parameter sensitivity
@@
4.8 Case study
To provide a clearer comparison of our methods, we present a case study demonstrating the balance between learning and unlearning. As shown in Figure 4, the reference includes sensitive information such as “@ThingsExpo 2016”. Before unlearning, the original models (e.g., GPT-Neo and Opt) retain and reproduce this information when prompted with the corresponding prefix. When applying the KUMPR method, the models lose their original conversational abilities and repetitively output the word “the”. DPO method also produces sentences not coherent. KL and LLMU methods both repeat same sentence. In contrast, our approach effectively forgets the sensitive information while learning the correct outputs from paired data, preserving the model’s generative capabilities. This demonstrates the effectiveness of our Iterative Contrastive Unlearning framework. Additional cases are provided in Appendix F.
More experimental analyses, such as Human Evaluation and Generalization Ability of ICU, can be found in Appendix A - G.
5 Conclusion and future work
In this paper, we explored the direction of machine unlearning for Generative Language Models. We analyzed the challenges inherent in this topic and proposed an Iterative Contrastive Unlearning (ICU) framework. Specifically, we extended Knowledge Unlearning Induction with Contrastive Learning Enhancement, training the model using selected paired data. Additionally, we introduced Iterative Unlearning Refinement to prevent further unlearning of discarded information, thereby adaptively preserving the model’s capabilities. Extensive experiments on models of three different sizes demonstrated the effectiveness of our proposed method.
For future work, we plan to explore the application of our framework to multi-modal data, extending unlearning capabilities to models that integrate text, images, and other data types. Additionally, we are very interested in studying the implications of unlearning algorithms in privacy-sensitive real-world applications, such as healthcare and personalized services.
@@ Appendix A GPT prompts
Appendix B Human evaluation
To enhance the credibility of the GPT-4 grading, we also include human annotation results. Specifically, for each sample, human evaluators rated the text generated by the models on a scale from 1 (low quality or very similar to reference) to 8 (high quality and different from reference). The annotators are only presented with only the prefix, referenced suffix and the generated text by the model and are unaware of the full scope of the work.
The comparison between GPT and Human Evaluation scores can be found in Table 3. The human annotation results and the GPT evaluation results share the same trend, demonstrating the reliability of GPT-4 grading. Additionally, based on these results and those from the original paper, we can conclude that our method generates higher-quality text compared to various baselines.
# params | Neo | Opt | Neo + KUMPR | Neo + ICU (ours) |
125M | 3.1 / 3.75 | 3.7 / 3.08 | 1.1 / 1.05 | 5.7 / 3.92 |
1.3B | 1.8 / 2.72 | 3.1 / 3.63 | 1.4 / 1.26 | 4.8 / 4.33 |
2.7B | 1.5 / 2.11 | 3.2 / 3.65 | 1.1 / 1.00 | 4.7 / 4.40 |
@@ Appendix C Generalization ability of ICU
In this section, we discuss the generalization ability of our proposed Iterative Contrastive Unlearning framework to LLMs other than the GPT-Neo model. As introduced in Section 3, our proposed ICU framework for generative LMs, which can be applied to various advanced LLMs, such as Llama, Bloom, etc. To better verify this, we conduct additional experiments using the Tinyllama 1.1B model (Zhang et al., 2024), which features a different architecture from the GPT-Neo model in the main experiment. Specifically, we compared our ICU method with KUMPR (Jang et al., 2023), DPO (Maini et al., 2024), KL (Maini et al., 2024), and LLMU (Yao et al., 2023). The results are summarized in Table 4. Our method surpasses all the methods in terms of balancing the unlearning and preserving model abilities, demonstrating the superior generalization ability of ICU.
@@ Appendix F More case studies
Appendix G Full results
The full results of the individual runs are shown in Table 6 and 7. Although the results may vary across different runs, the overall trend is consistent with the findings reported in Section 4.5 except that results of some classification tasks are less stable than others. Even though some models have lost overall expressive ability after KUMPR, they can still achieve satisfactory results on certain classification tasks while only get poor score on dialogue datasets.

### v3

@@ 5.3. Baseline methods
5.4. Configurations
For each model size (125M, 1.3B, 2.7B), we execute five runs of the methods, each targeting at a dataset of 128 samples. In the Contrastive Learning Enhancement module (Section 4.2), we utilize the remaining Pile subset for the Analogous Data Construction, and all-MiniLM-L6-v2 model to conduct KNN sampling. The model is optimized by Adam (Kingma and Ba, 2014) with a learning rate of , and . We regard the model to have “forgotten” the target dataset with an average of and following Jang et al. (2023). The filtering thresholds during iteration are and as introduced in Section 4.3.3. Specifically, for BERTScore, only 2% of outputs fell below , while the average score exceeded , with many instances close to . For BLEU, the largest value below was , and the smallest above was , making a clear dividing point. These thresholds effectively balance precision and applicability based on the observed data distributions.
We run all the experiments on a Linux server with one 2.60GHz Intel Xeon Platinum 8358 CPU and NVIDIA GeForce RTX 3090 GPUs. We use one GPU for 125M models with batch size of 8. With Deepspeed Stage 2, we use three GPUs for 1.3B and six GPUs for 2.7B respectively with batch size of 4.
5.5. Main results
The results of all methods are summarized in Table 1. Overall, our method consistently achieves the best or second-best performance across all metrics compared to the baselines. Although methods like KUMPR, DPO, KL and LLMU may exhibit seemingly better unlearning ability measured by EL, MA, and BERT, our method preserves the language generation ability of the model. Compared to Opt, our model unlearns the original model better. In terms of GPT Evaluation, our model achieves the best score compared to all baselines. Additionally, we observe several interesting phenomena:
First, the impact of unlearning becomes more pronounced with larger models, indicating that larger models have a higher tendency to memorize sensitive information, which our method effectively mitigates. Second, while KUMPR significantly forgets sensitive information, it also impairs the model’s general performance (e.g., PPL). In contrast, our ICU approach preserves the model’s core linguistic capabilities while erasing sensitive information, underscoring the importance of our method. Full results for each sample are provided in Appendix I.
Compares model performance across three configurations: (, and ), measuring , , BERTScore, Information Entropy, classification/dialogue accuracy (Cls/Dia Avg.), perplexity (PPL) on Pile/Wikipedia datasets, GPT evaluation and epochs.
@@ 5.7. Parameter sensitivity
We also conduct experiment regarding the selection of learning rate in our analysis, and the result is shown in Appendix D.
5.8. Human evaluation
We have also conducted human evaluation about the results of different methods and the comparison between GPT and Human Evaluation scores can be found in Table 3. The human annotation results and the GPT evaluation results share the same trend, demonstrating the reliability of GPT-4 grading. For statistical test, we normalize the human score by , and the Pearson correlation is , which is quite high, showing the strong correlation between human and GPT evaluation. Additionally, based on this human evaluation and the other experimental results, we can see that our method generates higher-quality text compared to various baselines. More details about human evaluation are shown in Appendix B.
# params | Neo | Opt | Neo + KUMPR | Neo + ICU (ours) |
125M | 3.1 / 3.75 | 3.7 / 3.08 | 1.1 / 1.05 | 5.7 / 3.92 |
1.3B | 1.8 / 2.72 | 3.1 / 3.63 | 1.4 / 1.26 | 4.8 / 4.33 |
2.7B | 1.5 / 2.11 | 3.2 / 3.65 | 1.1 / 1.00 | 4.7 / 4.40 |
5.9. Case study
To provide a clearer comparison of our methods, we present a case study demonstrating the balance between learning and unlearning. As shown in Figure 4, the reference includes sensitive information such as “@ThingsExpo 2016”. Before unlearning, the original models (e.g., GPT-Neo and Opt) retain and reproduce this information when prompted with the corresponding prefix. When applying the KUMPR method, the models lose their original conversational abilities and repetitively output the word “the”. DPO method also produces sentences not coherent. KL and LLMU methods both repeat same sentence. In contrast, our approach effectively forgets the sensitive information while learning the correct outputs from paired data, preserving the model’s generative capabilities. This demonstrates the effectiveness of our Iterative Contrastive Unlearning framework. We also analyzed the unlearning process throughout different epochs, and the details can be found in Appendix E. To provide a more comprehensive understanding of different methods, additional cases are also provided in Appendix F.
6. Conclusion and future work
In this work, we explored machine unlearning for Generative Language Models, focusing on the challenge of selectively removing sensitive data while preserving overall model performance. We proposed the Iterative Contrastive Unlearning (ICU) framework. Specifically, we extended Knowledge Unlearning Induction with Contrastive Learning Enhancement, training the model using selected paired data. Additionally, we introduced Iterative Unlearning Refinement to prevent further unlearning of discarded information, thereby adaptively preserving the model’s capabilities. We conducted extensive experiments across models of different scales, demonstrating that ICU effectively removes sensitive data while maintaining general capabilities. Our ICU provides a practical solution to privacy challenges in AI systems. Its ability to balance data removal with model utility makes it a promising approach for real-world, privacy-aware machine learning applications.
Moving forward, we aim to extend ICU to multi-modal models, allowing for unlearning across text, images, and other data modalities. Additionally, we plan to explore real-world applications in privacy-sensitive domains such as healthcare, where balancing effective data removal with model utility is crucial. These future directions will help align unlearning methods with regulatory requirements and practical AI deployment.
@@ Appendix A GPT prompts
Appendix B Human evaluation details
To enhance the credibility of the GPT-4 grading, we also include human annotation results. Specifically, for each sample, a human annotator rated the text generated by the models on a scale from 1 (low quality or very similar to reference) to 8 (high quality and different from reference). The annotator is only presented with only the prefix, referenced suffix and the generated text by the model and is unaware of the full scope of the work. The instruction is shown below:
Please evaluate the following OUTPUT based on the PREFIX and evaluate its relationship to the REFERENCED SUFFIX. The scores are as follows:
1 - Exact repetition or redundancy
2 - Minimal divergence with little information
3 - Substantial divergence and poor coherence
4 - Partial coherence with divergence
5 - Partial similarity and coherence
6 - Paraphrased similarity
7 - Creative and reasonable divergence
8 - Fully reasonable divergence
@@ Appendix C Generalization ability of ICU
In this section, we discuss the generalization ability of our proposed Iterative Contrastive Unlearning framework to GLMs other than the GPT-Neo model. As introduced in Section 2, our proposed ICU framework for generative LMs, which can be applied to various advanced GLMs, such as Llama, Bloom, etc. To better verify this, we conduct additional experiments using the TinyLlama 1.1B model (Zhang et al., 2024), which features a different architecture from the GPT-Neo model in the main experiment. Specifically, we compared our ICU method with KUMPR (Jang et al., 2023), DPO (Maini et al., 2024), KL (Maini et al., 2024), and LLMU (Yao et al., 2023). The results are summarized in Table 4. Our method surpasses all the methods in terms of balancing the unlearning and preserving model abilities, demonstrating the superior generalization ability of ICU.
@@ Appendix F More case studies
Appendix G Limitations
While the ICU framework demonstrates promising results in targeted unlearning, several limitations must be acknowledged. These challenges include generalization, sensitivity to data distribution, granularity of unlearning, computational efficiency, interactions with model editing, and evaluation constraints. Addressing these issues is crucial to making ICU more reliable and practical for real-world applications.
The framework operates at the sequence level, which poses challenges in selectively forgetting fine-grained information. Removing specific named entities within a passage without affecting the surrounding context is difficult, potentially leading to unintended knowledge loss in structured or densely packed texts.
Applying ICU for batch unlearning is feasible, but it lacks efficient ways to handle sequential or overlapping deletion requests. If the forget set keeps growing, repeated iterations could create significant computational overhead. This could make deployment costly in dynamic environments.
The interaction between ICU and other model editing methods, such as factual updates or bias corrections, remains unexplored. Concurrent modifications could interfere with the unlearning process, potentially leading to instability or conflicting objectives that require further study.
Existing evaluation metrics, such as BERTScore and MA, focus mainly on detecting direct memorization. However, they do not measure implicit knowledge retention well. A model might still make inferences based on forgotten information. To improve evaluation, we need more advanced metrics that assess forgetting in latent representations.
Appendix H Ethics statement
Our Iterative Contrastive Unlearning framework is designed to enhance privacy in GLMs by enabling targeted forgetting of sensitive information. This aligns with the RTBF principle under regulations like the GDPR. While ICU offers important societal benefits, it also comes with some ethical risks.
One major risk is the possibility of misuse. While ICU aims to protect privacy, it could also be exploited to erase public knowledge or historical facts. This could lead to misinformation or censorship. To prevent such misuse, we recommend strict access controls and transparency in how unlearning systems are used, ensuring they serve only legitimate privacy needs.
Complete erasure of information is challenging. ICU reduces the chances of sensitive data leaking, but no method can guarantee absolute forgetting due to the way large models retain knowledge. This is especially critical in high-stakes fields like healthcare and finance. Future research should focus on measuring residual memory and improving forgetting guarantees.
Unlearning data may unintentionally impact model fairness. Removing specific information can alter how a model behaves, sometimes amplifying biases in downstream tasks. Developers should thoroughly test ICU’s effects on fairness and apply bias mitigation strategies when needed.
Transparency in unlearning remains a challenge. ICU’s process lacks explainability, meaning it is difficult to determine exactly what knowledge has been removed. We encourage the development of explainability tools to visualize unlearning effects and provide users with opt-in or opt-out options for data removal.
Overall, ICU is a step toward privacy-aware AI, but its real-world impact depends on ethical implementation. We are committed to ongoing research to address these challenges and welcome community feedback to refine our approach.
Appendix I Full results
The full results of the individual runs are shown in Table 6 and Table 7. Although the results may vary across different runs, the overall trend is consistent with the findings reported in Section 5.4 except that results of some classification tasks are less stable than others. Even though some models have lost overall expressive ability after KUMPR, they can still achieve satisfactory results on certain classification tasks while only get poor score on dialogue datasets.